write an equation in slope-intercept form for the table of values:
X -3 -1 1 3
Y 7 5 3 1

Found 2 solutions by Alan3354, FrankM:
Answer by Alan3354(69443)

write an equation in slope-intercept form for the table of values:
X -3 -1 1 3
Y 7 5 3 1
--------
Pick 2 of the points
(1,3) & (3,1) will do
Slope m = diffy/diffx
m = -2/2 = -1
---
Use y = mx + b and any point to find b, the y-intercept
(3,1)
1 = (-1)*3 + b
b = 4
-------
y = -x + 4

Answer by FrankM(1040)

2 points define a line. (1,3) (3,1) is what I'll use.
slope-intercept is the form y=mx+b where m is the slope and b is the y intercept.
m=(y2-y1)/(x2-x1)= (1-3)/(3-1) = -2/2 = -1
with the slope, we substitute a point into this equation -
y-3=-1(x-1)
y-3=-1x+1 (distribute the -1 over the two terms)
y=-1x+4 (add 3 to both sides to simplify)
Now - you should double check that the equation works for the extra two pairs we didn't use. To be sure there were no mistakes along the way, and no typo in the question)